# Active solids: defect self-propulsion without flow

### v1

@@ Abstract
The self-propulsion of $+1/2$ topological defects is a hallmark of active nematic fluids, where the defects are advected by the flow field they themselves generate. In this paper we propose a minimal model for defect self-propulsion in a nematic active solid: a linear elastic medium with an embedded nematic texture that generates active stress and associated elastic strains. We show that such coupling gives rise to self-propelled $+1/2$ defects that move relative to the elastic medium by local remodeling of the nematic texture without advection. This mechanism is fundamentally different from the fluid case and can lead to unbinding of defect pairs and stabilization of $+1$ defects. Our findings might help explain how orientational order, of, for example, muscle fibers, is reconfigured during morphogenesis in solid-like tissues. The proposed mechanism may, for instance, control motility and merging of $+1/2$ defects, which play a crucial role in setting up the body axis during Hydra regeneration.
Topological defects are a common feature of orientational order and are found in many natural and synthetic systems. They are the organizing centers of orientational order fields and have been implicated in morphogenesis in vivo [1, 2, 3, 4] and in vitro [5], as well as in the organization of tissues [6, 7] and bacteria [8]. They have been widely studied in synthetic active fluids such as filament-motor mixtures, where nematic orientational order of cytoskeletal filaments directs force generation through molecular motors.
A hallmark of active nematic liquid crystals are self-propelled topological defects  [9]. In these fluids, defect motion is driven by advection of the nematic texture by the flow field generated by active stresses [10, 11]. Dynamic reconfiguration of nematic textures through defect creation, motion and fusion has been shown to play an important role during animal morphogenesis, for instance, during whole-body regeneration of the freshwater polyp Hydra [1, 4, 12, 3]. The tissues involved in these processes, however, have solid-like properties on the time scales of the defect dynamics [1, 2, 13], suggesting they should be described as active nematic solids rather than fluids. Rather than being advected by flow, defects may move relative to the surrounding elastic material as the nematic texture remodels locally [1, 2].
@@
In synthetic nematic elastomers, topological defects can be used to control shape morphing. In these materials internal forces can be generated through chemical or thermal activation of the elastomer, yielding nemato-elastic stresses that deform the system [14, 15, 16]. Similar ideas have recently been applied in the context of epithelial morphogenesis [17, 18, 19]. Theoretical studies of such shape-morphing solids have focused on quantifying the deformations induced by prescribed nematic textures [20, 21, 22].
@@
In this paper we examine topological defects in an active nematic elastomer, where forces are generated internally by active processes. We propose a new mechanism for defect motion through local melting of the nematic texture driven by strains induced by active stresses. This mechanism does not require flow of material, providing a natural framework for understanding the observed restructuring of the nematic texture in Hydra. We propose a minimal model for a linear elastic medium with an embedded nematic texture that generates active stress and in turn is coupled to elastic strain. We show that such coupling gives rise to self-propelled $+1/2$ defects that move relative to the elastic medium by local remodeling. If the self-propulsion is sufficiently large, $+1/2$ defects can escape the Coulomb-like attraction to a $-1/2$ defect. In confined systems where defects are enforced through boundary anchoring, the elasto-nematic coupling gives rise to rich dynamics, including orbiting $+1/2$ defects, and stabilized $+1$ defects that can spontaneously deform from aster/vortex configurations to spirals.
@@
where we have chosen coefficients such that $12Q_{ij}Q_{ij}=S^{2}=$ 1 for the uniform equilibrium when $\chi=\chi_{B}=0$. The coefficient $\chi$ accounts for strain alignment of nematic order along the direction of deviatoric strain $\tilde{u}_{ij}$, aligning them when $\chi>0$ and anti-aligning them for $\chi<0$. The coefficient $\chi_{B}$ describes the coupling of the nematic order to isotropic strain $u_{kk}$. We call this “strain ordering.” For $\chi_{B}>0$, nematic order decreases (“melts”) in regions of compressive strain ($u_{kk}<0$). Both strain alignment and strain ordering can either be of passive origin, or active, due to feedback-mediated regulation of the cytoskeletal organization in response to strain [25, 26, 2].
We consider adiabatic dynamics in a regime of quasi-static force balance where the nematic texture remodels on a time scale much slower than that of force equilibration through deformations. This assumption is justified in a tissue where elastic forces balance on short time scales (seconds to minutes) compared to the remodeling of the actomyosin fibers that form the nematic texture (hours). Force balance within the elastic sheet requires $\partial_{i}\sigma_{ij}=0$, where the total stress $\sigma_{ij}=\sigma^{\mathrm{el}}_{ij}+\sigma^{\mathrm{a}}_{ij}$ includes the passive elastic stress $\sigma^{\mathrm{el}}_{ij}=\frac{\delta\mathcal{F}}{\delta u_{ij}}$ and active isotropic and deviatoric stresses $\sigma^{\mathrm{a}}_{ij}=\frac{\alpha_{B}}{4}S^{2}\delta_{ij}+\alpha Q_{ij}$. Positive (negative) signs of $\alpha_{B}$ and $\alpha$ correspond to contractile (extensile) activity. Together, this gives
@@
where we have introduced the effective active stress strengths $\tilde{\alpha}_{B}=\alpha_{B}-\chi_{B}$ and $\tilde{\alpha}=\alpha-\chi$ that capture the total isotropic and deviatoric stresses exerted by the nematic. We neglect passive stresses due to the nematic molecular field which are higher order in gradients than the active nematic stresses. Throughout the manuscript, we set $K=\mu=B/2=1$. Unless specified otherwise, we set $\chi_{B}=-2$, $\tilde{\alpha}_{B}=0$.
@@
The effective active stresses induce elastic stresses which can relax through remodeling of the nematic texture. Geometrically, the induced elastic stresses can be understood as a consequence of metric incompatibility (intrinsic curvature) as is briefly discussed in Appendix A. As the texture remodels, the active stresses will, however, lead to further incompatibility. In other words, activity makes the coupling between elastic strain and nematic order non-reciprocal, thus allowing for the emergence of dynamic steady states [27, 28, 29], as shown below.
Figure 1: (a) Illustration of elasto-nematic couplings considered in our model. Black sticks illustrate nematic directors. Red and blue arrows show stress and strain respectively. (b) Gradient of isotropic strain across a $+1/2$ defect, causing the defect to self-propel (green arrow) through “strain ordering” in the tail (see Movie 1). (c) Steady-state defect position $r_{0}^{*}$, where propulsion force $\bm{f}^{+\frac{1}{2}}$ and Coulomb repulsion $\bm{f}_{C}$ due to anchoring at the boundary balance (see inset), as a function of active stress; numerics (circles) and analytic estimate (green line). (Parameters: $R=16,\tilde{\alpha}=1,\chi=0$.)
Self-propelled $+1/2$ defects. Only $+1/2$ defects have polar symmetry which is necessary for self-propulsion [30]. To investigate propulsion mechanisms in our model, consider a single $+1/2$ defect with its tail pointing along the direction $\hat{\mathbf{p}}$. Since we study an isolated defect, we can use rotational and translational invariance to set $\hat{\mathbf{p}}=\hat{\mathbf{x}}$, the positive $x$-direction and center the defect at the origin:
@@
where $(r,\theta)$ denote polar coordinates. Equation (4) describes the “unperturbed” equilibrium texture of a $+1/2$ defect with $\chi=\chi_{B}=0$. The magnitude of order $S_{1/2}(r)$ vanishes at the defect core $r=0$ and is $1$ for $r\gg\sqrt{K}$ with a transition zone (core size) of width $\sim\sqrt{K}$  [^1] An approximate solution for $S_{1/2}(r)$ can be found using the Pad‘e approximant [34] (see SI).
@@
which for $\mathbf{Q}=\mathbf{Q}_{+\frac{1}{2}}$ has the approximate solution $u^{+\frac{1}{2}}_{kk}(\mathbf{r})\approx-\frac{\tilde{\alpha}S_{1/2}(r)}{2(B+\mu)}\frac{x}{r}$ (see SI). For contractile active stress ($\tilde{\alpha}>0$), as expected for instance in a tissue like a Hydra’s ectoderm [32], this strain field is tensile in front of the defect and compressive behind it [see Fig. 1(d)]. In our model, this strain gradient will lead to a free energy density gradient across the defect through the strain-ordering coupling $\chi_{B}$. Thus, remodeling of the nematic texture, Eq. (3), will propel the defect in the direction that reduces this free energy. As the defect moves, the strain field generated by active stresses shifts along with it such that there will always be a strain gradient across the defect core, sustaining the effective force acting on the defect. This propulsion mechanism is markedly different from the fluid case, where the defect is advected by the velocity field it actively induces [10, 11].
To estimate the defect propulsion force, we formulate Eq. (3) variationally and make the ansatz $\mathbf{Q}(\mathbf{r},t)=\mathbf{Q}_{+\frac{1}{2}}[\mathbf{r}-\mathbf{r}_{0}(t)]$, which yields (see Appendix B for details)
@@
Importantly, the direction of propulsion of the $+1/2$ is determined by both the sign of the active stress $\alpha$ and that of the coupling parameter $\chi_{B}$. In particular, this allows a system with contractile activity to have a defect move towards its head. This is in contrast to active fluids, where the sign of activity alone controls the direction of propulsion such that $+1/2$ defects move towards their tail for contractile activity.
To numerically validate the analytic estimate for $\bm{f}_{\chi_{B}}^{+\frac{1}{2}}$, we simulate a $+1/2$ defect in disk with suitable anchoring boundary conditions. Boundary anchoring leads an effective force $\bm{f}_{\mathrm{ext}}=-\frac{\pi}{2}K\mathbf{r}_{0}/(R^{2}-r_{0}^{2})$ that pushes the defect towards the disk center and can be calculated by placing a $+1/2$ image charge at $(R^{2}/r_{0}^{2})\mathbf{r}_{0}$. Equation (9) gives a good prediction for the position where the defect stalls ($\bm{f}_{\chi_{B}}^{+\frac{1}{2}}+\bm{f}_{\mathrm{ext}}=0$); see Fig. 1(c).
In the complementary case of purely isotropic induced stress ($\tilde{\alpha}=0$, $\tilde{\alpha}_{B}\neq 0$), the isotropic strain is azimuthally symmetric and, hence, $\bm{f}_{\chi}^{+\frac{1}{2}}$ vanishes. Due to azimuthal symmetry, the deviatoric strain induced by isotropic stress has principal axes pointing along the radial and azimuthal directions. This implies that the strain alignment contribution $\chi\tilde{u}_{ij}Q_{ij}$ in the free energy has opposite signs in the head and the tail of the defect, as illustrated in Fig. 4(a) in Appendix C. The resulting defect-propulsion force $\bm{f}_{\chi}^{+\frac{1}{2}}\approx-\frac{\pi}{16}\frac{\tilde{\alpha}_{B}\chi\sqrt{K}}{B+\mu}\hat{\mathbf{p}}$ is derived in Appendix C.
@@
Defect-pair unbinding. In a passive nematic liquid crystal, defects of opposite charge attract through a Coulomb-like force with strength $\pi K/(2d)$ [33], where $d$ is the defect separation. We therefore expect that a $+1/2$ defect that is oriented away from the $-1/2$ defect can escape its attraction beyond the critical separation $d^{*}=-\frac{\pi\sqrt{K}\,(B+\mu)}{0.4\,\tilde{\alpha}\chi_{B}}\sim K/|\mathbf{v}_{0}|$, analogously to active fluids [10, 34]. To to test this prediction, we simulated ±$\pm 1/2$ defect pairs in a rectangular geometry as shown in Fig. 2(a) (see Movie 2). Indeed for strong enough active stress $\alpha$ and large enough initial separation $d$, the defect separation increases with time and the $+1/2$ defect escapes the $-1/2$ [see Fig. 2(b)].
@@
Interaction of $+1/2$ defects in a confined domain. Biological tissue is often geometrically confined which can enforce the presence of topological defects through boundary anchoring [5] or through topological constraints on surfaces with non-zero Euler characteristic (e.g. spheroids [1]). To study the interaction of two $+1/2$ defects in a confined geometry, we consider a disk-shaped domain with perpendicular nematic anchoring along the boundary which enforces a net $+1$ topological charge. For simplicity, we set $\tilde{\alpha}_{B}=0$ throughout this section.
@@
Figure 3: (a,b) Phase diagrams quantifying the steady-state defect separation (a) and “twist angle” (b) for two $+1/2$ defects in a disk geometry with normal boundary anchoring of the nematic. The dotted gray lines in (a,b) bound the region where $+1$ defects are stable (see SI). The white dot-dashed line in (a) shows the analytically predicted stability threshold [see Eq. (26)]. In the region which is black in (a), initially separated $+1/2$ defects merge to form a $+1$ defect. Below this is a bistable region, where $+1/2$ defects either repel or merge into a $+1$ depending on their initial distance (see SI). The red dashed lines show the linear stability prediction for spiral formation (see SI). (c) Representative snapshots for parameter combinations labeled (i–iv) in the phase diagram (see Movie 3). In the spiral regime, $+1/2$ defects orbit around the origin [indicated by green arrows in (i)]. (Domain radius $R=16$; free boundary conditions: $\sigma_{ri}|_{r=R}=0$.)
@@
Figures 3(a) and 3(b) show the numerical phase diagram in the $(\tilde{\alpha},\chi)$-plane where we characterize the configurations in terms of the defect separation $d$, Fig. 3(a), and the twist angle $\psi$, Fig. 3(b), measuring the orientation of the $+1/2$ polarization (head–tail axis) relative to the radial direction. Movie 3 and representative snapshots in Fig. 3(c) show the four distinct types of behaviors found in simulations: (i) a twisted nematic texture that spontaneously breaks chiral symmetry with two orbiting $+1/2$ defects; a stable $+1$ defect in the center of the domain (ii, iii); and (iv) a configuration where the $+1$ breaks into two $+1/2$s pushed to the boundary. For large negative $\chi$, we observe bistability (hysteresis) where the steady state configuration depends on the initial condition (see SI).
@@
The transition to twisted nematic textures for $\tilde{\alpha}\chi<0$ can be understood in terms of geometric incompatibility: for a $+1$ defect with phase $\psi$ [the director angle relative to the displacement vector from the defect core; see inset of Fig. 3(b)] the nematogen-induced stress generates intrinsic Gaussian curvature of magnitude $\sim\cos(2\psi)$ concentrated at the core [21, 22] (see Appendix D). Therefore, in a flat geometry, there will be residual stresses unless $\psi=\pm\pi/4$. Thus, starting from an achiral state $\psi=0,\pi/2$, relaxation of elastic energy will spontaneously break chiral symmetry to $\psi=\pm\pi/4$ within a boundary layer of width $\sim\sqrt{K}$. The onset of spiral formation as a function of domain size can be estimated by a linear stability analysis (see SI), which is in good agreement with simulations [Fig. 3(b)].
A $+1$ defect forms in the center of the disk when the self-propulsion of the two $+1/2$s is strong enough to push them closer than the core radius. The active stress required to stabilize a $+1$ defect in the center can be estimated from energetic arguments, as shown in Appendix D. The resulting estimate yields the white dashed line in Fig. 3(a) which agrees well with numerics.
@@
Role of substrate traction. The ectoderm of Hydra is separated by the endoderm by a layer of extracellular matrix that serves as a substrate to the tissue. The role of a rigid substrate can be incorporated by introducing a spring-like traction force $-ku_{i}$ which modifies force balance to read $\partial_{j}\sigma_{ij}=ku_{i}$. Long-ranged elastic interactions then become screened, with a screening length $\ell_{\mathrm{el}}\sim\sqrt{(B+\mu)/k}$. Such a screening plays a role analogous to that of the viscous screening length in active fluids [35] and eliminates system-size dependence arising form the long-range nature of elastic forces (e.g., in the stability condition for $+1$ defects Eq. (26), see Appendix D).
@@
Importantly, traction provides an additional mechanism for feedback between mechanics and nematic order, as it allows a term $\lambda u_{i}u_{j}Q_{ij}$ in the free energy Eq. (1) describing alignment of $Q_{ij}$ to the displacement field. This may represent, for instance, alignment of actin fibers along the direction of traction force. In the limit of strong coupling to the substrate, we can neglect elastic stresses and solve the force balance equation for the displacement field as $u_{i}=\frac{\alpha}{k}\partial_{j}Q_{ij}$. Alignment with displacement then gives a contribution $(\lambda\alpha^{2}/k)\partial_{k}Q_{ik}\partial_{l}Q_{jl}$ to the dynamics of $Q_{ij}$. Interestingly, this term generates a defect propulsion with direction independent of the sign of activity and determined entirely by the sign of the alignment interaction. For parallel alignment of nematic order and traction, $+1/2$ defects will always move towards their head.

### v2

# Active Solids: Topological Defect Self-Propulsion Without Flow

@@ Abstract
The self-propulsion of +1/2+1/2 topological defects is a hallmark of active nematic fluids, where the defects are advected by the flow field they themselves generate. In this paper we propose a minimal model for defect self-propulsion in a nematic active solid: a linear elastic medium with an embedded nematic texture that generates active stress and associated elastic strains. We show that such coupling gives rise to self-propelled +1/2+1/2 defects that move relative to the elastic medium by local remodeling of the nematic texture without advection. This mechanism is fundamentally different from the fluid case and can lead to unbinding of defect pairs and stabilization of +1+1 defects. Our findings might help explain how orientational order, of, for example, muscle fibers, is reconfigured during morphogenesis in solid-like tissues. The proposed mechanism may, for instance, control motility and merging of +1/2+1/2 defects, which play a crucial role in setting up the body axis during Hydra regeneration.
@@
Topological defects are found in many natural and synthetic systems. They are the organizing centers of orientational order fields and have been implicated in morphogenesis in vivo [1, 2, 3, 4] and in vitro [5, 6, 7], as well as in the organization of tissues [8, 9] and bacterial colonies [10]. In synthetic nematic elastomers, topological defects can be used to control shape morphing. In these materials internal forces can be generated through chemical or thermal activation of the elastomer, yielding nemato-elastic stresses that deform the system [11, 12, 13]. Theoretical studies of such shape-morphing solids have focused on quantifying the deformations induced by prescribed (“frozen-in”) nematic textures [14, 15, 16, 17]. On the other hand, dynamic reconfiguration of nematic textures through defect creation, motion and fusion has been shown to play an important role in morphogenesis, as studied for instance, during whole-body regeneration of the freshwater polyp Hydra [1, 4, 18, 19].
Self-propelled defect dynamics has been studied extensively in active nematic liquid crystals [20]. In these fluids, active stresses generate flows which in turn drive defect motion through advection [21, 22]. By contrast, defects in solid-like epithelial tissues have been observed to move relative to the surrounding elastic material as the nematic texture remodels locally [1, 2]. In such tissues, cells adhere to one another strongly and their rearrangements happen much slower than the remodeling of intracellular filaments [23, 1, 24, 7], calling for a description as active nematic solids rather than fluids.
@@
Here, we investigate the dynamics of topological defects in an active nematic elastomer, where forces are generated internally by active processes. We propose a new mechanism for defect motion through local melting of the nematic texture driven by strains induced by active stresses. This mechanism does not require flow of material, providing a natural framework for understanding the observed restructuring of the nematic texture in Hydra. We propose a minimal model for a linear elastic medium with an embedded nematic texture that generates active stress and in turn is coupled to elastic strain. We show that such coupling gives rise to self-propelled +1/2+1/2 defects that move relative to the elastic medium by local remodeling. If the self-propulsion is sufficiently large, +1/2+1/2 defects can escape the Coulomb-like attraction to a 1/2-1/2 defect. In confined systems where defects are enforced through boundary anchoring, the elasto-nematic coupling gives rise to rich dynamics, including orbiting +1/2+1/2 defects, and stabilized +1+1 defects that can spontaneously deform from aster/vortex configurations to spirals.
@@
where we have chosen coefficients such that $2Q_{ij}Q_{ij}=S^{2}=1$ for the uniform equilibrium when $\chi=\chi_{B}=0$. The coefficient $\chi$ accounts for strain alignment of nematic order along the direction of deviatoric strain $\tilde{u}_{ij}$, aligning them when $\chi>0$ and anti-aligning them for $\chi<0$. The strain alignment term can be derived by linearizing the Warner–Terentjev model for nematic elastomers [11, 27]. It leads to a soft mode since the elastomer can accommodate shear through reorientation of the nematic. The coefficient $\chi_{B}$ describes the coupling of the nematic order to isotropic strain $u_{kk}$. We call this “strain ordering.” For $\chi_{B}>0$, nematic order decreases (“melts”) in regions of compressive strain ($u_{kk}<0$). To our knowledge, this term has not been previously considered (with the exception of Ref. [28] where only defect-free systems are studied), since most nematic elastomer theories assume incompressibility. In the context of biological tissue, both strain alignment and strain ordering can result from feedback-mediated regulation of the cytoskeletal organization in response to strain [29, 30, 2]. In this case $\mathcal{F}$ should be understood as an effective free energy.
We consider adiabatic dynamics in a regime of quasi-static force balance where the nematic texture remodels on a time scale much slower than that of force equilibration through deformations. This assumption is justified in a tissue where forces balance on short time scales (seconds to minutes) compared to the remodeling of the actomyosin fibers that form the nematic texture (hours). Force balance within the elastic sheet requires $\partial_{i}\sigma_{ij}=0$, where the total stress $\sigma_{ij}=\sigma^{\mathrm{el}}_{ij}+\sigma^{\mathrm{a}}_{ij}$ includes the passive elastic stress $\sigma^{\mathrm{el}}_{ij}=\frac{\delta\mathcal{F}}{\delta u_{ij}}$ and active isotropic and deviatoric stresses $\sigma^{\mathrm{a}}_{ij}=\frac{\alpha_{B}}{4}S^{2}\delta_{ij}+\alpha Q_{ij}$. Positive (negative) signs of $\alpha_{B}$ and $\alpha$ correspond to contractile (extensile) activity. Together, this gives
@@
where we have introduced the effective active stress strengths $\tilde{\alpha}_{B}=\alpha_{B}-\chi_{B}$ and $\tilde{\alpha}=\alpha-\chi$ that capture the total isotropic and deviatoric stresses exerted by the nematic  [^1] If strain alignment and strain ordering result from biological feedback loops rather than passive (thermodynamic) mechanisms, $\mathcal{F}$ is only an effective free energy and the alignment terms do not contribute to the stress $\chi$ and $\chi_{B}$ in the stress. In this case, $\alpha=\tilde{\alpha}$, $\alpha_{B}=\tilde{\alpha}_{B}$.. We neglect passive stresses due to the nematic molecular field which are higher order in gradients than the active nematic stresses. Throughout the manuscript, we set $K=\mu=B/2=1$. Unless specified otherwise, we set $\chi_{B}=-2$, $\tilde{\alpha}_{B}=0$.
@@
The effective active stresses induce elastic stresses which can (partially) relax through remodeling of the nematic texture. This is akin to the soft elasticity of nematic elastomers that accommodate external deformation by reorientation of the nematic [11]. Geometrically, the induced elastic stresses can be understood as a consequence of metric incompatibility (intrinsic curvature) as is briefly discussed in Appendix A. As the texture remodels, the active stresses will, however, lead to further incompatibility. In other words, activity makes the coupling between elastic strain and nematic order non-reciprocal, thus allowing for the emergence of dynamic steady states [32, 33, 34], as shown below.
Figure 1: (a) Illustration of the elasto-nematic couplings considered in our model. Black sticks illustrate nematic directors. Red and blue arrows show stress and strain respectively. (b) Gradient of isotropic strain across a +1/2+1/2 defect, causing the defect to self-propel (green arrow) through “strain ordering” in the tail (see Movie 1). (c) Steady-state defect position $r_{0}^{*}$, where propulsion force $\bm{f}^{+\frac{1}{2}}$ and Coulomb repulsion $\bm{f}_{C}$ due to anchoring at the boundary balance (see inset), as a function of active stress; numerics (circles) and analytic estimate (green line). (Parameters: $R=16,\tilde{\alpha}=1,\chi=0$.)
Self-propelled +1/2+1/2 defects. Only +1/2+1/2 defects have polar symmetry which is necessary for self-propulsion [35]. To investigate propulsion mechanisms in our model, consider a single +1/2+1/2 defect with its tail pointing along the direction $\hat{\mathbf{p}}$. Since we study an isolated defect, we can use rotational and translational invariance to set $\hat{\mathbf{p}}=\hat{\mathbf{x}}$, the positive $x$-direction and center the defect at the origin:
@@
where $(r,\theta)$ denote polar coordinates. Equation (4) describes the “unperturbed” equilibrium texture of a +1/2+1/2 defect with $\chi=\chi_{B}=0$. The magnitude of order $S_{1/2}(r)$ vanishes at the defect core $r=0$ and is $1$ for $r\gg\sqrt{K}$ with a transition zone (core size) of width $\sim\sqrt{K}$  [^2] An approximate solution for $S_{1/2}(r)$ can be found using the Pad‘e approximant [40] (see SI).
@@
which for $\mathbf{Q}=\mathbf{Q}_{+\frac{1}{2}}$ has the approximate solution $u^{+\frac{1}{2}}_{kk}(\mathbf{r})\approx-\frac{\tilde{\alpha}S_{1/2}(r)}{2(B+\mu)}\frac{x}{r}$ (see SI). For contractile active stress ($\tilde{\alpha}>0$), as found for instance in a tissue like a Hydra’s ectoderm [37], this strain field is tensile in front of the defect and compressive behind it [Fig. 1(b)]. In our model, this strain gradient leads to a free energy density gradient across the defect through the strain-ordering coupling $\chi_{B}$. Thus, remodeling of the nematic texture, Eq. (3), propels the defect in the direction that reduces this free energy. As the defect moves, the strain field generated by active stresses shifts along with it such that there will always be a strain gradient across the defect core, sustaining the effective force acting on the defect. This propulsion mechanism is markedly different from the fluid case, where the defect is advected by the velocity field it actively induces [21, 22].
To estimate the defect propulsion force, we make the variational ansatz $\mathbf{Q}(\mathbf{r},t)=\mathbf{Q}_{+\frac{1}{2}}[\mathbf{r}-\mathbf{r}_{0}(t)]$, which yields (see Appendix B for details)
@@
Importantly, the direction of propulsion of the +1/2+1/2 defect is determined by both the sign of the active stress $\alpha$ and that of the coupling parameter $\chi_{B}$. In particular, this allows a system with contractile activity to have a defect move towards its head when $\chi_{B}$ is negative, i.e. when extensile strain reduces nematic order. This is in contrast to active fluids, where the sign of activity alone controls the direction of propulsion such that +1/2+1/2 defects move towards their tail for contractile activity.
To numerically validate the analytic estimate for $\bm{f}_{\chi_{B}}^{+\frac{1}{2}}$, we simulate a +1/2+1/2 defect in a disk with suitable anchoring boundary conditions. Boundary anchoring leads to an effective force $\bm{f}_{\mathrm{ext}}=-\frac{\pi}{2}K\mathbf{r}_{0}/(R^{2}-r_{0}^{2})$ that pushes the defect towards the disk center and can be calculated by placing a +1/2+1/2 image charge at $(R^{2}/r_{0}^{2})\mathbf{r}_{0}$. Equation (9) gives a good prediction for the position where the defect stalls ($\bm{f}_{\chi_{B}}^{+\frac{1}{2}}+\bm{f}_{\mathrm{ext}}=0$); see Fig. 1(c). Deviations are due to the variational ansatz that neglects distortions of the nematic texture [38] and to approximations made when estimating the elastic strain.
In the complementary case of purely isotropic induced stress ($\tilde{\alpha}=0$, $\tilde{\alpha}_{B}\neq 0$), the isotropic strain is azimuthally symmetric and, hence, $\bm{f}_{\chi}^{+\frac{1}{2}}$ vanishes. Due to azimuthal symmetry, the deviatoric strain induced by isotropic stress has principal axes pointing along the radial and azimuthal directions. This implies that the strain alignment contribution $\chi\tilde{u}_{ij}Q_{ij}$ in the free energy has opposite signs in the head and the tail of the defect [Fig. 4(a)]. The resulting defect-propulsion force $\bm{f}_{\chi}^{+\frac{1}{2}}\approx-\frac{\pi}{16}\frac{\tilde{\alpha}_{B}\chi\sqrt{K}}{B+\mu}\hat{\mathbf{p}}$ is derived in Appendix C.
@@
Defect-pair unbinding. In a passive nematic liquid crystal, defects of opposite charge attract through a Coulomb-like force with strength $\pi K/(2d)$ [39], where $d$ is the defect separation. We therefore expect that a +1/2+1/2 defect that is oriented away from the 1/2-1/2 defect can escape its attraction beyond the critical separation $d^{*}=-\frac{\pi\sqrt{K}\,(B+\mu)}{0.4\,\tilde{\alpha}\chi_{B}}\sim K/|\mathbf{v}_{0}|$, analogously to active fluids [21, 40]. To test this prediction, we simulated $\pm 1/2$ defect pairs in a rectangular geometry [Fig. 2(a); Movie 2]. Indeed for strong enough active stress $\alpha$ and large enough initial separation $d$, the defect separation increases with time and the +1/2+1/2 defect escapes the 1/2-1/2 [Fig. 2(b)].
@@
Interaction of +1/2+1/2 defects in a confined domain. Biological tissue is often geometrically confined which can enforce the presence of topological defects through boundary anchoring [5] or through topological constraints on surfaces with non-zero Euler characteristic (e.g. spheroids [1]). To study the interaction of two +1/2+1/2 defects in a confined geometry, we consider a disk-shaped domain with perpendicular nematic anchoring along the boundary enforcing a net +1+1 topological charge.
@@
Figure 3: (a,b) Phase diagrams quantifying the steady-state defect separation (a) and “twist angle” (b) for two +1/2+1/2 defects in a disk geometry with normal boundary anchoring of the nematic. The regions in the top left and bottom right corner where $\psi>0$ and $d>0$, correspond to orbiting defect pairs. The dotted gray lines in (a,b) bound the region where +1+1 defects are stable (see SI). The white dot-dashed line in (a) shows the analytically predicted stability threshold [see Eq. (26)]. In the region which is black in (a), initially separated +1/2+1/2 defects merge to form a +1+1 defect. Above this is a bistable region, where +1/2+1/2 defects either repel or merge into a +1+1 depending on their initial distance (see SI). The red dashed lines show the linear stability prediction for spiral formation (see SI). (c) Representative snapshots for parameter combinations labeled (i–iv) in the phase diagram (see Movie 3). In the spiral regime, +1/2+1/2 defects orbit around the origin [indicated by green arrows in (i)]. (Domain radius $R=16$; free boundary conditions: $\sigma_{ri}|_{r=R}=0$.)
@@
Figures 3(a) and 3(b) show the numerical phase diagram in the $(\tilde{\alpha},\chi)$-plane where we characterize the configurations in terms of the defect separation $d$ and the twist angle $\psi$ measuring the orientation of the +1/2+1/2 polarization (head–tail axis) relative to the radial direction. Movie 3 and representative snapshots in Fig. 3(c) show the four distinct types of behaviors found in simulations: (i) a twisted nematic texture that spontaneously breaks chiral symmetry with two orbiting +1/2+1/2 defects; a stable +1+1 defect in the center of the domain (ii, iii); and (iv) a configuration where the +1+1 breaks into two +1/2+1/2s pushed to the boundary. For large $\chi$, we observe bistability (hysteresis) where the steady state configuration depends on the initial condition (see SI).
@@
The transition to twisted nematic textures for $\tilde{\alpha}\chi<0$ can be understood in terms of geometric incompatibility: For a +1+1 defect with phase $\psi$ (the director angle relative to the polar angle from the defect core) the nematogen-induced stress generates intrinsic Gaussian curvature of magnitude $\sim\cos(2\psi)$ concentrated at the core [15, 16] (see Appendix D). Therefore, in a flat geometry, there will be residual stresses unless $\psi=\pm\pi/4$. Starting from an achiral state $\psi=0,\pi/2$, relaxation of elastic energy will spontaneously break chiral symmetry to $\psi=\pm\pi/4$ within a boundary layer of width $\sim\sqrt{K}$. A linear stability analysis taking into account this boundary layer [see SI; dashed red lines in Fig. 3(a,b)] yields good agreement with simulations. In a large part of the parameter space, the +1+1 charge is resolved into a pair of +1/2+1/2 defects. Due to the spiral-like texture, the defects’ head-tail axes are at an angle to the radial direction, such that their self-propulsion drives cyclic motion [Fig. 3(c-i)].
A +1+1 defect forms in the center of the disk when the inward pointing radial component of the defects’ self-propulsion force is strong enough to push them closer than the core radius. The active stress required to stabilize a +1+1 defect in the center can be estimated from energetic arguments (Appendix D), yielding good agreement with numerics [Fig. 3(a)]. Notably, this “active stabilization” becomes more effective in large domains, in contrast to passive +1+1 defects that are stabilized by confinement of size similar to the nematic coherence length [41]. Alternative mechanisms for active stabilization of +1+1 defects have recently been studied in models with shape-tension coupling [42, 43].
@@
Role of substrate traction. Tissues studied in vitro are often attached to a rigid substrate [5, 8, 9] which can be incorporated by introducing a spring-like traction force $ku_{i}$ so that force balance reads $\partial_{j}\sigma_{ij}=ku_{i}$. Long-ranged elastic interactions then become screened, with a screening length $\ell_{\mathrm{el}}\sim\sqrt{(B+\mu)/k}$, analogously to viscous screening in active fluids [44]. Screening eliminates system-size dependence arising form the long-range nature of elastic forces (e.g., in the stability condition for +1+1 defects Eq. (26), see Appendix D).
@@
Importantly, traction provides an additional mechanism for feedback between mechanics and nematic order, as it allows a term $\lambda u_{i}u_{j}Q_{ij}$ in the free energy Eq. (1) describing alignment of $Q_{ij}$ to the displacement field. This may represent, for instance, alignment of actin fibers along the direction of traction force. In the limit of strong coupling to the substrate, we can neglect elastic stresses and solve force balance for the displacement field as $u_{i}\approx\frac{\tilde{\alpha}}{k}\partial_{j}Q_{ij}$ to leading order. Alignment with displacement then gives a contribution $(\lambda\tilde{\alpha}^{2}/k)\partial_{k}Q_{ik}\partial_{l}Q_{jl}$ to the dynamics of $Q_{ij}$. For parallel alignment of nematic order and displacement/traction, this causes +1/2+1/2 defects to move towards their head, independent of the sign of activity. In a purely passive system, the defect propulsion through this mechanism will be transient until distortion of the nematic texture leads to arrest of motion, as has recently been shown in nematic fluids [38].